# Can Large Language Models Address Open-Target Stance Detection?

### v1

@@ Abstract
Stance detection (SD) assesses a text’s position towards a target, typically labeled as "favor," "against," or "neutral." We introduce Open-Target Stance Detection (OTSD), where targets are neither seen during training nor provided as input. Evaluating Large Language Models (LLMs) like GPT-3.5, Llama 3, and Mistral, we compare their performance with the Target-Stance Extraction (TSE) approach, which has the advantage of using predefined targets. LLMs perform better than TSE in target generation when the real target is explicitly and not explicitly mentioned in the text. For stance detection, LLMs perform better in explicit scenarios but fail in non-explicit ones. [^1] We will release the code and the dataset.
@@
Abu Ubaida Akash  and Ahmed Fahmy  and Amine Trabelsi {akaa2803, mosa2801, amine.trabelsi}@usherbrooke.ca Université de Sherbrooke
@@ 1 Introduction
Stance detection (SD) is the task of determining the position of a text or person towards a certain target, often split into “favor” and “against” with an optional third split being “neutral”, “none”, or “neither” (Sutter et al., 2024). The target word can be mentioned in the text explicitly, or sometimes the idea of the target word can be conveyed indirectly (Küçük and Can, 2020). In Zero-shot Stance Detection (ZSSD), the model predicts the stance for targets that it has not seen during the training phase. This task is essential because collecting training data for every possible topic or target is impractical (Allaway et al., 2021). While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022a; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022; Liang et al., 2022b; Li and Yuan, 2022), most studies assume that the target is known or manually identified and given as input, a situation that is uncommon in real-world applications. In practice, the target is unknown, and it is not necessarily mentioned explicitly in the text. Additionally, target annotation is a tedious and expensive task in SD (Küçük and Can, 2020; Sutter et al., 2024).
@@
In this paper, we focus on a different yet challenging task we refer to as Open-Target Stance Detection (OTSD). In this task, the target is neither seen during training (zero-shot) nor provided as input to the model. In OTSD, one important challenge lies in identifying the target from the text rather than having it provided as input. The target may be mentioned explicitly or not. The stance is predicted with respect to the produced target.
There has been one notable attempt to partially address this real-world setting of SD. Li et al. (2023a) generate targets from the input text and later detect stance based on these targets. However, before starting the SD task, they map these generated targets to a predefined list of golden targets (details in §2) to ensure exact wording and facilitate evaluation. They called their approach Target-Stance Extraction in the Wild (TSE). This approach does not fully align with our OTSD setting, which assumes no input target information is given during the whole process, and is impractical due to the need for a comprehensive list of all possible targets. Moreover, our experiment reveals a performance gap in Li et al. (2023a)’s approach when used on text with explicit or non-explicit target mentions. This aspect was not highlighted by the authors and warrants further investigation in the context of OTSD.
@@
In this work, we present and explore the OTSD task, focusing on its two main steps: Target Generation and Stance Detection (see Eq. 2). As a zero-shot task for both target and stance, Large Language Models (LLMs) appear well-suited for this challenge. Therefore, we examine the performance of LLMs and compare it to TSE, the primary existing work by Li et al. (2023a), which benefits from using a predefined list of targets during the process (see Eq. 1). Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—specifically GPT-3.5 (Brown et al., 2020), Llama 3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023)—perform in open target generation compared to the TSE approach when the real target is explicitly or not explicitly mentioned in the text for stance classification, and how do they compare to each other? [RQ#2] How do the same LLMs perform compared to TSE in stance detection (in the context of OTSD) for both explicit and non-explicit cases, and how do they compare to each other?
Our contributions consist of introducing the task of OTSD and conducting experiments and answer these research questions.
@@ 2 Target-Stance Extraction (TSE)
In TSE, Li et al. (2023a) demonstrate target extraction and stance detection tasks in both in-target and zero-shot settings. For our comparison, we only consider their zero-shot setting. In TSE zero-shot, they first generate the target using keyphrase generation models, map the generated targets (multiple targets per sample) to a predefined list of golden targets to find the most similar one, and finally use it to detect the stance of the text toward the target (Eq. 1). The best match from their keyphrase model’s output, marked as TSE-BestGen, is used for our comparison. The best mapped target (TSE-mapped) is considered the predicted target. For example, given the text "Embracing different faiths teaches us empathy and respect. Let’s celebrate our diverse beliefs and foster unity. #Faith #Unity #Respect", TSE primarily generates targets such as Peace, Religious diversity, Respect, etc.. From a predefined list of possible targets (e.g., Face Mask, Atheism, Donald Trump, etc.), Atheism matches most closely with Religious diversity and is considered as the final predicted target. In OTSD, we aim at directly generating Religious diversity or a very semantically similar phrasing, avoiding the constraints of a predefined list. In the second step, they detect the stance toward their generated target.
@@ 3.2 Approach
To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs. We adopt the "Task Definition" prompting strategy for the OTSD task, as it has been shown by Cruickshank and Ng (2023) to outperform other strategies—such as "Task-only", "Context Analyze", "Context Question", "Zero-shot CoT", and "CoDA"—particularly for smaller models (7B-8B parameters) across most datasets. Given that this is an open-target task, we aim to use a straightforward prompting approach to maintain fairness and generalizability; we do not tailor the prompt for specific targets or include demonstrations or in-context examples, as this remains a zero-shot task.
Since LLMs are sensitive to the length of generated text, we limit the target length to ensure a fair comparison with TSE in terms of target generation quality. To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets. Based on our empirical study, we set the maximum target length to 5 for GPT-3.5 and 4 for Llama and Mistral to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ). Finally, we employ two different prompting approaches for target generation (TG) and stance detection (SD).
TG+SD (Two-Step Approach)
In this approach, we first generate the target and then sequentially detect the stance of the text toward the generated target. By directing the LLM to focus on one task at a time, this method may reduce the cognitive load on the model.
TG&SD (Single-Step Joint Approach)
This approach combines target generation and stance detection into a single step using a unified prompt. By jointly generating the target and stance, the model can develop a more holistic understanding of the relationship between the text, target, and stance.
The prompts and their design justifications can be found in Appendix A.
@@ 4 Experiments
In the OTSD task, we compare our approaches with TSE in both TG and SD. To ensure sensitivity and robustness, we run our experiments three times independently and report the average results  [^2] Score variations are negligible, up to four decimal places.. Since OTSD is a generative task, we primarily compare our results with TSE-BestGen, despite its reliance on a predefined list of golden targets. Additionally, we consider TSE-mapped as the current best score, regardless of task type. Details about our hardware settings can be found in Appendix D.
@@ 4.1 Dataset and Model
We use the same dataset as TSE’s zero-shot setup for a fair comparison, consisting of a total of 3,000 samples considering three stance classes i.e., Favor, Against, and None. This includes 500 samples each from 6 different targets (see Appendix E). Additionally, we categorize the dataset into explicit and non-explicit samples. This categorization involves removing stop words and special characters from both texts and targets, lemmatizing all words, and subsequently comparing the text with the target. A sample is labeled as explicit if any target words are found in the corresponding text; otherwise, it is labeled as non-explicit. This process results in 1,804 explicit samples and 1,196 non-explicit samples (in Appendix E). It is important to note that because we compare our approach with TSE zero-shot, and they do not provided the code or sufficient details for reproducing the generation or selection of candidate targets, we are unable to use other stance detection datasets for benchmarking.
In our focused contribution, we utilize both proprietary and non-proprietary LLMs, including GPT-3.5 (Brown et al., 2020), Llama-3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023), to address our two research questions (§1). [^3] Although we do not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs is financially constraining for us. Model descriptions can be found in Appendix  B.
@@ 4.2 Evaluation Methods
To measure the generated target quality in OTSD task, we consider the following metrics:
BTSD: We fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE (fine-tuning details in Appendix F). The SD score (F1-macro) is used as the evaluator for the generated targets. To assess the effectiveness of this metric, we experiment with different levels of target quality by modifying or removing words from the golden targets, substituting them with incorrect targets, and selecting random words from the vocabulary as targets. We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (Appendix C, Table 3). Therefore, the F1 score on stance detection serves as a reliable proxy for evaluating target generation quality in our dataset.
Human Evaluation: To evaluate the quality of each generated target and to further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit, with 16.67% from each stance label). Three independent domain experts are asked to assess the relevance of the generated targets by evaluating its relevance to the golden target in the tweet. Each annotator is provided with an individual file, ensuring that their evaluations remain independent of one another. The full agreement details and score distribution are included in Appendix I and J, respectively.
SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related. Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings. Detailed steps can be found in Appendix G.
@@
TG+SD TG&SD
Model SemSim BTSD HEval SemSim BTSD HEval
Explicit
TSE-M 0.96 36.63 0.716 - - -
TSE-B 0.86 35.8 0.338 - - -
GPT-3.5 0.87 38.43 0.546 0.87 39.6 0.663
Llama-3 0.88 38.31 0.54 0.87 37.75 0.558
Mistral 0.85 38.47 0.586 0.87 39.02 0.65
Non-explicit
TSE-M 0.9 30.56 0.391 - - -
TSE-B 0.82 29.32 0.215 - - -
GPT-3.5 0.84 33.1 0.419 0.83 31.32 0.405
Llama-3 0.85 33.06 0.447 0.83 33.81 0.49
Mistral 0.83 30.18 0.386 0.83 34.12 0.495
Table 1: Comparison of TG performance between TSE and LLMs across various evaluation metrics. HEval represents human evaluation. TSE-M and TSE-B correspond to TSE-mapped and TSE-BestGen, respectively. The highest score for each metric across all models is highlighted in bold, while for BTSD and HEval, the better results between the two approaches are underlined.
@@ 5 Result Analysis
5.1 Target Generation
Following RQ#1, Table 1 shows that for the explicit case, Mistral achieves the highest BTSD score in the TG+SD setting, while GPT-3.5 leads in TG&SD. Overall, LLMs outperform TSE-BestGen in the BTSD metric and produce more relevant targets according to HEval scores. Although TSE-mapped has the highest SemSim and HEval scores, its BTSD score is lower due to generating either a perfect match or something entirely different (as illustrated in Figure 1). In contrast, the LLMs generate targets that are either highly related or partially related to the golden ones, with fewer irrelevant targets compared to TSE-mapped, leading to higher BTSD scores despite lower HEval. For non-explicit cases, LLMs outperform TSE-BestGen in both BTSD and HEval, with Llama-3 generating the most relevant targets (as indicated by the HEval score), even surpassing TSE-mapped. In the explicit case, TG&SD generally outperforms TG+SD across most LLMs, except for Llama, which performs better with TG+SD in BTSD. In non-explicit cases, GPT excels with TG+SD, while the other LLMs perform better with TG&SD. Among LLMs, GPT performs best overall in the explicit case, while Mistral leads in the non-explicit case. A manual inspection (on GPT-3.5’s 30 incorrect predictions) investigating the lower score in non-explicit cases reveals that the model struggles without sufficient context, often due to the way the TSE test set is constructed. This lack of context, such as implicit hints on the target or the context of other messages in the thread, causes the model to generate targets based solely on the provided text (example in Appendix H.1), leading to inaccuracies. Table 1 also confirms that higher HEval scores correlate with higher BTSD scores, validating BTSD as a reliable proxy for evaluating target quality.
Model TG+SD TG&SD
Explicit
TSE-mapped 38.1 -
TSE-BestGen 37.81 -
GPT-3.5 42.68 47.61
Llama-3 43.13 42.1
Mistral 46.15 46.32
Non-explicit
TSE-mapped 32 -
TSE-BestGen 31 -
GPT-3.5 32.66 33.94
Llama-3 33.36 31.9
Mistral 33.61 30.47
Table 2: Comparison of SD performance between TSE and LLMs, measured by F1-macro (in %). The highest scores across all models are highlighted in bold, with the best scores between the two approaches underlined.
5.2 Stance Detection
To address RQ#2, Table 2 shows that in stance detection, LLMs without a predefined list outperform TSE, which uses a predefined list, in both approaches (TG+SD and TG&SD) for explicit cases. In non-explicit cases, the scores are comparable in the TG+SD setting, but higher scores are observed for LLMs in the TG&SD setting. This is because, in explicit cases, LLMs generate better-quality targets than TSE-BestGen, leading to improved stance detection as evidenced in Table 1. For non-explicit cases, the holistic view provided by the TG&SD setting makes it easier for LLMs to infer stance from the text. Note that although LLM generated targets are better quality than TSE-BestGen, their stance labels can still be different from golden labels, resulting in low F1 scores. Our manual analysis revealed that these errors mainly arise from insufficient context in the input text, both explicit and non-explicit (see Appendix H.2 for examples). In both approaches, LLMs achieve much higher SD scores in explicit cases compared to non-explicit ones. Again, this is due to the incorrectly generated targets in the non-explicit case revealed by manual inspection. Incorrect targets correspond in a high percentage to incorrect stance labels. Most of targets were either completely incorrect or, in a few cases, contradictory to the real target, altering the stance.

### v3

@@ Abstract
Stance detection (SD) identifies a text’s position towards a target, typically labeled as favor, against, or none. We introduce Open-Target Stance Detection (OTSD), the most realistic task where targets are neither seen during training nor provided as input. We evaluate Large Language Models (LLMs) GPT-4o, GPT-3.5, Llama-3, and Mistral, comparing their performance to the only existing work, Target-Stance Extraction (TSE), which benefits from predefined targets. Unlike TSE, OTSD removes the dependency of a predefined list, making target generation and evaluation more challenging. We also provide a metric for evaluating target quality that correlates well with human judgment. Our experiments reveal that LLMs outperform TSE in target generation when the real target is explicitly and not explicitly mentioned in the text. Likewise, for stance detection, LLMs excel in explicit cases with comparable performance in non-explicit in general. [^1] Dataset and code are available here: anonymous.4open.science/r/opentarget-5521
[…]
@@
Abu Ubaida Akash  Ahmed Fahmy  Amine Trabelsi Université de Sherbrooke {akaa2803,mosa2801,amine.trabelsi}@usherbrooke.ca
@@ 1 Introduction
Stance detection (SD) aims to determine the position of a text or person towards a certain target, typically categorized as “favor”, “against”, or “none”. The target can be mentioned explicitly in the text, or sometimes the idea of the target can be conveyed indirectly (Küçük and Can, 2020). In Zero-shot Stance Detection (ZSSD), a model predicts stances for targets it has not seen during training, which is crucial since collecting training data for every potential target is impractical (Allaway et al., 2021). While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022), most studies assume that the target is known or manually identified and given as input, a rare scenario in real-world applications where the target is often unknown or not explicitly mentioned. Additionally, target annotation is an expensive task in SD (Küçük and Can, 2020).
@@
In this paper, we focus on a different yet challenging task we refer to as Open-Target Stance Detection (OTSD). In this task, the target is neither seen during training (zero-shot) nor provided as input to the model. In OTSD, one important challenge lies in identifying the target from the text rather than having it provided as input. The target may be mentioned explicitly or not and the stance is predicted with respect to the produced target.
@@
There has been one notable attempt to partially address this real-world setting of SD. Li et al. (2023a) generate targets from the input text and later detect stance based on these targets. However, in target generation, they map the generated targets to a predefined list of golden targets (details in §2) to ensure exact wording and facilitate evaluation. They call their approach TSE, which does not fully align with our OTSD setting. In OTSD we assume no input target information is given during the whole process, making it more practical than TSE, which requires a comprehensive list of all possible targets. Moreover, our experiment reveals a performance gap in Li et al. (2023a)’s approach when used on text with explicit or non-explicit target mentions, a factor not addressed by the authors that warrants further investigation in the context of OTSD. In this work, we present and explore the OTSD task, focusing on its two main steps: Target Generation (TG) and Stance Detection (SD) (Eq. 2). As a zero-shot task for both target and stance, Large Language Models (LLMs) appear well-suited for this challenge. Therefore, we examine the performance of LLMs and compare it to the primary existing work, TSE, which benefits from using a predefined list of targets during the process (Eq. 1). Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—specifically GPT-4o (Achiam et al., 2023), GPT-3.5 (Brown et al., 2020), Llama 3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023)—perform in open TG compared to the TSE when the real target is explicitly or non-explicitly mentioned in the text for stance classification, and how do they compare to each other? [RQ#2] How do the same LLMs perform compared to TSE in SD (in the context of OTSD) for both explicit and non-explicit cases, and how do they compare to each other? Our contributions consist of introducing the task of OTSD, providing a target evaluation metric, and conducting experiments answering the research questions.
@@ 2 Target-Stance Extraction (TSE)
In TSE, Li et al. (2023a) demonstrate target extraction and stance detection tasks in both in-target and zero-shot settings. For our comparison, we only consider their zero-shot setting. In TSE zero-shot, they first generate targets using keyphrase generation models, then map these targets to a predefined list of golden targets to find the closest match, and finally use this match to detect the stance of the text (Eq. 1). The best match from the keyphrase model, marked as TSE-BestGen, and the best mapped target (TSE-mapped) is considered as the predicted target. For example, given the text "Embracing different faiths teaches … diverse beliefs and foster unity.", TSE primarily generates targets such as Peace, Religious diversity, Respect, etc.. From a predefined list of possible targets (e.g., Face Mask, Atheism, Donald Trump, etc.), Atheism matches most closely with Religious diversity and is considered as the final predicted target. In OTSD, we aim to generate Religious diversity or a closely related phrase directly, bypassing the predefined list. The stance is then detected toward this generated target.
@@ 3.2 Approach
To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs. We adopt the "Task Definition" prompting strategy, as it has been shown by Cruickshank and Ng (2023) to outperform other strategies—such as "Task-only", "Context Analyze", "Context Question", "Zero-shot CoT", and "CoDA"—particularly for smaller models (7B-8B parameters) across most datasets. Since this is an open-target task, we use a straightforward prompting approach to ensure fairness and generalizability, avoiding tailored prompts for specific targets, or in-context examples, aligning with the zero-shot nature of this task. We employ two different prompting approaches for TG and SD. The prompts and their design justifications can be found in Appendix A. TG+SD (Two-Step Approach) In this approach, we first generate the target and then sequentially detect the stance of the text toward the generated target. By directing the LLM to focus on one task at a time, this method may reduce the cognitive load on the model. TG&SD (Single-Step Joint Approach) This approach combines TG and SD into a single step using a unified prompt. By jointly generating the target and stance, the model can develop a more holistic understanding of the relationship between the text, target, and stance.
@@ 4 Experiments
In OTSD task, we compare our approaches with TSE in both TG and SD. To ensure sensitivity and robustness, we run our experiments three times independently and report the average results  [^2] Score variations are negligible, up to four decimal places.. Since OTSD is a generative task, we primarily compare our results with TSE-BestGen, despite its reliance on a predefined list of golden targets. Additionally, we consider TSE-mapped as the current best score, regardless of task type. Details about our hardware settings can be found in Appendix D.
@@ 4.1 Dataset and Model
To ensure a fair comparison between TSE and LLM models, we use the same dataset as TSE’s zero-shot setup. VAST (Allaway and McKeown, 2020) is used to compare LLMs across different settings. Both TSE and VAST datasets contain three stance classes (Favor, Against, None), and contain 3,000 and 5,100 samples, respectively, in our study. VAST, originally a multi-target dataset, is converted to a single-target format (conversion process in Appendix E) to match the scope of our study. TSE has 500 samples for each of 6 targets (see Appendix F), while VAST includes 2,145 unique targets. We classify samples as explicit or non-explicit by removing stop words and special characters, lemmatizing all words, and checking for target words in the text. This results in 1,804 explicit and 1,196 non-explicit samples from TSE, and 3,120 explicit and 1,980 non-explicit samples from VAST (see Appendix F). Note that, due to the lack of code and details from TSE for reproducing target generation and selection, we cannot use other SD datasets for benchmarking. In our focused contribution, we use both proprietary and non-proprietary LLMs, including GPT-3.5, GPT-4o, Llama-3, and Mistral to address our two research questions (§1).
@@ 4.2 Evaluation Methods
BTSD: To measure the generated target quality, we fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE (fine-tuning details in Appendix G). The SD score (F1-macro) is used as the evaluator for the generated targets. To assess the effectiveness of this metric, we experiment with different levels of target quality by modifying or removing words from the golden targets, substituting them with incorrect targets, and selecting random words from the vocabulary as targets. We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (see Appendix C). More importantly, BTSD score shows strong tau ($\tau$) correlation (KENDALL, 1938) of $0.85$ and $0.79$ for TSE and VAST, respectively, with human judgments gathered during experimentation (see the section below). Therefore, the BTSD serves as a reliable proxy for evaluating target generation quality.
Human Evaluation: To evaluate the relevance of the generated targets and further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit) from both datasets, maintaining an equal ratio of the stance labels. Three annotators were asked to assess the relevance of the generated targets to the golden ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related), following the guidelines provided in Appendix J. To measure inter-annotator agreement among the three raters, we compute Krippendorff’s $\alpha$ (Krippendorff, 2011), and Fleiss’s $\kappa$ (Fleiss, 1971) for each model, as shown in Appendix Table 5. We observe a solid overall agreement of $\alpha$=0.76 and $\kappa$=0.664.
SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related. Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings. Detailed steps can be found in Appendix H with human judgement correlation of $\tau$=$0.57$ for TSE, and $\tau$=$0.59$ for VAST.
@@
TG+SD TG&SD TG+SD TG&SD
Model SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC
Explicit Explicit
[…]
TSE-B 0.86 35.8 0.338 37.81 - - - - - - - - - - - -
[…]
Non-explicit Non-explicit
[…]
TSE-B 0.82 29.32 0.215 31 - - - - - - - - - - - -
[…]
Table 1: Comparison of TG and SD performance between TSE and LLMs using various metrics: TG is evaluated with SS (SemSim), BTSD (%), and HE (Human Evaluation) scores, while SC (%) measures SD performance. TSE-M and TSE-B represent TSE-mapped and TSE-BestGen, respectively. Metric-wise best results within a specific setting are underlined, and better results of a model across TG+SD and TG&SD for each dataset are in bold.
@@ 5 Result Analysis
5.1 Target Generation (TG)
As shown in Table 1, both human evaluation (HE) and BTSD indicate that LLMs generate higher-quality targets in explicit cases compared to TSE-BestGen, the best model for generating targets not matched to a predefined list. Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower. This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix K). In contrast, LLMs mostly generate targets that are highly or partially related to golden targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
When the target is not explicitly mentioned, LLMs produce higher-quality targets than both TSE models. However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 1), with the best-performing LLM getting the HE score slightly lower than $0.5$ (refers to partially related target) on average. A manual inspection of GPT-3.5’s 30 incorrect predictions suggests that the lower score in non-explicit cases is due to insufficient context, often resulting from how the TSE test set is constructed. The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (see example in Appendix I.1).
When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa). We note that GPT-4o consistently produces targets of equal or higher quality than other tested models across the settings, including explicit and non-explicit target mentions, TG+SD, and TG&SD approaches, and across the datasets. Among the tested open LLMs, Mistral generates higher quality targets than Llama-3 in most settings. Comparing between the approaches, most models perform better in TG&SD across both datasets, for explicit and non-explicit cases.
5.2 Stance Detection (SD)
As shown in Table 1, according to the SC (stance classification), LLMs outperform both TSE models in detecting stances for explicit cases, while performing comparably with TSE models in non-explicit cases (except GPT-4o).
In general, LLMs perform better in SD when the target is explicitly mentioned in the text compared to when it is not. Our manual analysis shows that in non-explicit cases, the input text lacks sufficient context about the target, leading to lower SC scores.
In general, the experimented proprietary models (GPT-3.5 and GPT-4o) outperform the open ones in TG&SD in both the explicit and non-explicit cases across the datasets. In some cases of TG+SD setting, though, the open LLMs achieve comparatively higher scores. Comparing approaches, LLMs tend to perform better when prompted with the holistic TG&SD, though their TG+SD results remain competitive, specifically in non-explicit cases across both datasets.
Although the quality of the generated targets correlates positively with the SC score in explicit cases (TSE $\tau$=$0.71$, VAST $\tau$=$0.29$), this correlation is negative in non-explicit cases (TSE $\tau$=$-0.14$, VAST $\tau$=$-0.29$). We manually investigate the negative correlation (lower SC with higher quality target) in non-explicit cases. Indeed, LLMs such GPT-4o may generate highly relevant targets that are closely related to the golden targets but nearly ‘antithetical’ or opposite, leading to a stance reversal in comparison to the golden stance. For example, when the golden target is "permits to carry guns", GPT-4o generates "gun control in universities", while GPT-3.5 generates "guns". Although GPT-4o’s target is more closely related to the golden target, it represents the opposite stance. On the other hand, GPT-3.5’s target, though broader, less accurate and less related, aligns with the stance of golden target. This highlights the need for a coherence measure in OTSD that accounts for the target-stance alignment.